OpenStudy (anonymous):

the question says to solve algebraically and graphically. x+1-2√(x+4)=0 I did the algebraic part but my answer could not be solved

OpenStudy (anonymous):

If you're looking for x: x+1 = 2sqrt(x+4) Square both sides x^2 + 2x + 1 = 4(x+4) ====> x^2 + 2x + 1 = 4x + 16 Move everything to the right x^2 + 2x - 4x + 1 -16 = 0 X^2 -2x -15 = 0 Find your roots (x - 5)(x + 3) = 0 So x is either 5 or -3

OpenStudy (anonymous):

thank you so much best response so far and detailed

OpenStudy (anonymous):

to solve graphically do I make an xy char and plug in numbers

OpenStudy (anonymous):

Yup and just see where y =0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do I plug the numbers into x^2-2x-15